OpenStudy (linn99123):

A square garden has an area of 103 square feet. To the nearest tenth, what is the length of one side of the garden? A. 51.5 feet B. 25.8 feet C. 10.1 feet D. 7.2 feet

OpenStudy (31356):

B

OpenStudy (bibby):

\[x*x = x^2 = 103\]

OpenStudy (russano):

All you have to do is square each answer choice (multiply it by itself) to find the answer.

OpenStudy (bibby):

\[\sqrt{x^2}=x\]

OpenStudy (anonymous):

the answer is 51.5